Small quantitles of axygen can be prepared in the laboratory by heating potassum chlorate, KCios(9 The equation for the reactión is Calculate how many grams of Og) can be produced from heating 66.8 grams of KCIOs(s) Number 0 g o
media%2F7d6%2F7d6cdbf4-b13c-4b4e-96bd-08
Answer #1

mass of KClO3 = 66.8 g

molar mass of KClO3 = 122.55 g/mol

mol of KClO3 = (mass)/(molar mass)

= 66.8/122.55

= 0.5451 mol

According to balanced equation

mol of O2 formed = (3/2)* moles of KClO3

= (3/2)*0.5451

= 0.8176 mol

mass of O2 = number of mol * molar mass

= 0.8176*32

= 26.2 g

Answer: 26.2 g

I am allowed to answer only 1 question at a time